UNIT CODE: BAF2202 UNIT TITLE: MANAGEMENT ACCOUNTING I                           

  1. Assume that a certain process has an 95% learning curve effect and the first unit took 4000hrs to produce.

Required:

  1. Compute the number of hours required to produce the first 34 units
  2. Compute the number of hours required to produce 34th unit.

iii) Assume that the wage rate is Kshs.100 per hour, compute the Labor cost of producing the last 16 units

  1. Differentiate between management accounting and financial accounting
  2. Highlight the assumptions of cost volume profit analysis
  3. Describe any five cost estimation methods applied in management accounting
Answer #1
A
A learning curve is geometric with the general form Y = aX^b.
Y = cumulative average time per unit or batch.
a = time taken to produce initial quantity.
X = the cumulative units of production or, if in batches, the cumulative number of batches.
b = the learning index or coefficient, which is calculated as : log learning curve percentage ÷ log 2. So b for an 95 percent curve would be log 0.95 ÷ log 2 = – 0.074
(i) the number of hours required to produce the first 34 units.
Y34 = 4000 * (34)^-0.074
Y34 = 4000 * 0.7703
Y34 = 3081.2 Hours
Cumulative Number of hours to produce the first 34 units = 3081.2 * 34 units
Cumulative Number of hours to produce the first 34 units = 104760.80 Hours
(ii) the number of hours required to produce the 34th unit.
the number of hours required to produce the first 33 units.
Y33 = 4000 * (33)^-0.074
Y33 = 4000 * 0.7720
Y33 = 3088.0 Hours
Cumulative Number of hours to produce the first 33 units = 3088.0 * 33 units
Cumulative Number of hours to produce the first 33 units = 101904 Hours
Now,
the number of hours required to produce 34th unit is
   = 3081.2 * 34 Units - 3088.0 * 33 Units
   = 104760.80 - 101904
2856.8
   = 2856.8 Hours
the number of hours required to produce the 34th unit is 2856.80 Hours.
(iii) the Labor cost of producing the last 16 units.
Cumulative Number of hours to produce the 34 units (point (i)) = 104760.80 Hours
Cummulative Number of hours to produce the first 18 units = 4000 * (18)^-0.074 * 18 = 58132.80 Hours.
Cumulative Number of hours to produce the last 16 Units = 104760.80 - 58132.80 = 46628 Hours
Labor cost of producing the last 16 units = 46628 Hours * Kshs.100 per hour = 4,662,800.
B. Difference between Management Accounting & Financial Accounting
(i) Managerial (management) accounting provides information that is useful in running a company by internal users, whereas, the financial accounting is the preparation of financial reports for the analysis by the external users interested in knowing the financial position of the company. This is the key difference between financial accounting and management accounting.
(ii) Financial accounting is required by law while management accounting is not. Standards are available for both.
(iii) Financial accounting is much broader than Management accounting, Because, management accounting may be concerned with particular products or cost centers/departments, while, Financial accounting covers the whole organization.
(iv) Financial accounting is required to be in accordance with specific Standards, Formats prescribed as applicable, whereas, in Management Accounting, no specific format is designed.
(v) Financial accounting focuses on the past, whereas, in Management Accounting, the focus is on the Future.
C. The assumptions of cost volume profit analysis.
(i) Costs can be categorized as variable or fixed.
(ii) No change in Sale Price, Variable Cost, and Fixed Cost. All three are constant.
(iii) The only driver is the Number of Units, i.e. Volume.
(iv) No Semi Variable/Semi Fixed Cost.
(v) All units produced are sold during the period.
D. Five cost estimation methods applied in management accounting
Different cost estimation techniques are used by different industries as applicable to them. The following are the generally accepted cost estimation methods.
1. High-low method
The high-Low method is a quick and easy approach to estimate cost. The high-low method uses historical information from several reporting periods to estimate the costs. The variable cost is measured by subtracting the cost of the lowest activity from the cost of the highest activity and dividing the resultant amount from the difference of units of highest activity and the units of the lowest activity.
2. Scattergraph method
The Scattergraph method is used to segregate mixed costs and is more accurate than the above-mentioned high-low method because the scattergraph method considers all data points, not just the highest and lowest levels of activity. All the data points are plotted on the graph. Then a line is fitted dividing the points, with half the points above the line, and half the points below the line. This is called a regression line. The next steps help estimate the Fixed Cost and Variable Cost per unit. The results are put into the equation: Y = F + Vx.
3. Account analysis
The account analysis approach is the most basic approach for estimating fixed and variable costs. The experienced employee of the concerned department reviews and makes analysis of the appropriate accounts and determines the nature(Fixed/Variable) of the cost. This is one of the way the company estimates its costs.
4. Budgeting Estimate
Budgeting or also known as Preliminary Cost Estimation. Under this method, major historical data is used to estimate the cost and changes in cots behavior. There is various other methods under this head Budgeting Estimate. The main purpose of this type of estimation is to generate a particular list of expenditures/costs. this is done within the actual planning stage of the project.
5. Regression Analysis
This Technique for estimating costs describes the relationship between variables. It helps understand how the value of a dependent variable changes when one independent variable is variable while another is held constant. Used majorly for forecasting future data. This method is similar to the Scattergraph Approach for Cost Estimation.